RLC frequency limits Last week you studied the driven RC circuit and the driven RL circuit. Consider the driven series RLC circuit. Its impedance is Z=R+j(ωL−1/(ωC)) .

(a) For very high frequencies the driven RLC circuit behaves like an RL circuit (the effect of the capacitor becomes negligible). In this case does the current lead or lag the applied voltage in phase?

A. lags B. leads C. neither

What value in degrees does |ϕ| approach as ω→∞ ?

(b) For very low frequencies the driven RLC circuit behaves like an RC circuit (the effect of the inductor becomes negligible). In this case does the current lead or lag the applied voltage in phase?

A. leads B. lags C. neither

What value in degrees does |ϕ| approach as ω→∞ ?

(c) At resonance the RLC series circuit acts like a pure resistance R. In this case does the current lead or lag the applied voltage in phase?

A. leads B. lags C. neither

What angle in degrees is |ϕ| at resonance?
Answer #1

A.

In inductive (RL) circuit,current lags the voltage.

tan =( L/ R)

At tending to infinity, tan = infinite

= 90 o

B.

In capacitive circuit (RC) , current leads the voltage.

tan= 1/​​​​CR

as omega tends to infinity, tan = 0

or, = 0o

C.

At resonance, both current and voltage are in phase together.

option c is correct.

At rsonareso, XC=X​​​​​​L

1/C = L

tan = XC - X​​​​​​L/R = 0

so, = 0o